OpenStudy (anonymous):

Integrate 1/((x^2 -5x +6)^.5). I am having trouble with this and would a appreciate any help.

OpenStudy (anonymous):

\[\int\limits\frac{ 1 }{ \sqrt{(x^2 -5x +6)} } \]

OpenStudy (anonymous):

It isn't a trig sub, but i am not sure what else to do.

OpenStudy (anonymous):

just a second

OpenStudy (anonymous):

Im realy sorry but ill find help for you is that ok say yes if your there

OpenStudy (anonymous):

@Jadeishere @Jamierox4ev3r @jswagger @JayBaybie

OpenStudy (anonymous):

@Warped_909

OpenStudy (anonymous):

I am here, sorry afk'